ADVANCES IN CW CAVITY RING-DOWN SPECTROSOCPY

\underline{M. D. LEVENSON}, M. D. Levenson Consulting, 19868 Bonnie Ridge Way, Saratoga, CA 95070.; B. A. PALDUS, T. G. SPENCE, R. N. ZARE, Dept. of Chemistry, Stanford University, Stanford, CA 94305-5080.; B. WILLKE, C. C. HARB, R. L. BYER, Ginzton Laboratories, Stanford University, Stanford, CA 94305..

Stable continuous wave (cw) Nd:YAG and external cavity diode lasers, servo-locked to ring cavities enhance the performance of cavity ring-down spectroscopy (CRDS). Low-noise cw light that is not coupled into the long-lived cavity mode can be used as the local oscillator in heterodyne detection. This technique reduces the impact of electronic noise and facilitates shot-noise-limited sensitivity even with vanishingly small ring-down signals. Two variations have been demonstrated: frequency-switched and polarization-selective optical heterodyne detection. Spectra of CO_{2} at 1.06 μm were obtained signal-to-noise ratios above 40 dB, limited only by digitization error.